# Why collective behaviours self-organise to criticality: A primer on information-theoretic and thermodynamic utility measures

### v1

@@ I Introduction
This principle encapsulates the intrinsic utility of self-organising collective behaviour, elucidating why some systems gravitate towards criticality. Our discussion will begin with a background on criticality and phase transitions provided in Section II. Section III outlines the notations and essential technical preliminaries. Section IV provides an overview of established intrinsic utility approaches, while Section V describes thermodynamic efficiency defined at the intersection of information theory and thermodynamics. A common example using the two-dimensional Ising model is presented in Section VI to compare the self-organising behaviours driven by different intrinsic utilities. Section VII offers a more in-depth discussion on the principle of super-efficiency, and Section VIII summarises the findings and presents the final conclusion. This study offers insights into the distinct characteristics of collective behaviour derived within each framework, and emphasises how the principle of super-efficiency captures the intrinsic utility for collective behaviour at the critical regime.
@@ IV.2.1 Predictive information
$\begin{split}\mathcal{I}:&=\quad I(S_{past};S_{future})\\ &=\underbrace{H(S_{future})}_{\text{Richness of future states}}-\underbrace{H(S_{future}|S_{past})}_{\text{Unpredictability of future}}\end{split}$ (8)
@@ IV.3 Beyond intrinsic motivation
Figure 9: The diagram illustrates interaction between elements in the active inference framework. Solid lines represent influences between components. Dash lines represent directed influence from sensory to internal or from action to external, which correspond to the two stages of active inference. Figure is adapted from [49].
@@
However, the quantity called in this approach “free energy” is different from the thermodynamic free energy. In active inference, it is instead the variational free energy formulated in terms of information theoretic quantities, relating to the Bayesian inference process [45, 49]. Informally, anything that can be represented in the form:
@@
While equation (18) implies a connection between minimizing free energy and reducing energy cost under Landauer’s principle [78], the relationship is not explicitly formulated as a ratio of informational gain to energetic costs.
@@ V Thermodynamic efficiency
At this stage we point out that the three information-theoretic approaches reviewed in the previous section do not explicitly account for the corresponding energy costs. Thermodynamic efficiency, on the other hand, takes into consideration both the benefits and the associated costs of maintaining order within the system. Before presenting a formal definition of thermodynamic efficiency, it is important to differentiate between thermal and thermodynamic efficiency.
@@ V.1 Thermal vs thermodynamic efficiency
Let us consider a system undergoing a non-ideal process in which it receives energy and performs useful work. Not all the received energy is converted into work; some is inevitably lost as heat, which does not contribute to work output (Figure 10). Thermal efficiency measures the system’s efficiency of converting energy to work, and is defined as the ratio of useful work output to total energy input, both measured in joules, rendering it a dimensionless quantity. In a non-ideal process, the second law of thermodynamics implies that this ratio is less than one.
In contrast, thermodynamic efficiency assesses the conversion of work into the system order, measured during a quasi-static change in the underlying control parameter. It pertains to systems involving interactions among multiple components, and considers the benefit of increasing order within a collective system against the thermodynamic cost incurred. A system may transition from a disordered to an ordered state by altering a control parameter according to a specific protocol. Thermodynamic efficiency evaluates how efficiently the system converts the carried out work into order, at each specific value of the control parameter (Figure 11). It is quantified as the ratio of the reduction in the system’s configuration entropy (predictability gain) to the generalised work performed during the control parameter adjustment, expressed in units of bits per joule (subject to the unit of the Boltzmann constant $k_{B}$, e.g., see the expression for entropy, defined in the context of thermodynamics in Equation (21)):
@@
Figure 11: Thermodynamic efficiency for a system. It is defined as the ratio between the increase in order (measured in bits) and change in the generalised work carried out to generate the order (measured in joules). This quantity has a unit of bits per joule.
@@ V.3 Fisher information
At this stage we turn our attention to the Fisher Information which is related to Gibbs free energy. The Fisher information measures the amount of information that an observable random variable carries about an unknown parameter $\theta$ which may be influencing the probability of observations. In order words, Fisher information quantifies the sensitivity of observations to the change of parameter $\theta$. Mathematically, Fisher information is defined as the variance of the score function, where the score is the derivative of the log-likelihood function with respect to $\theta$ [61]:
@@ VI.1 The 2D Ising model
The 2D Ising model offers a simplified representation of ferromagnetism in statistical mechanics. It models a collection of sites that can each exhibit either an up-spin or down-spin configuration, while interacting with their neighbours to create a complex aggregate dynamic. The 2D Ising model is particularly relevant to our comparison due to its characteristic phase transition in the collective dynamics.
@@
In this example, the dynamics of the Ising model is interpreted from a perception-action loop perspective: each site acts as an agent that senses the energy of its neighbourhood and “decides” whether to flip its spin or maintain its current state. The agency of each site is determined by the choice of the coupling parameter $J$. With a high $J$ value a site is more prone to align with its neighbours, and vice versa. This parameter governs the strength of the site’s response to the neighbourhood’s energy landscape, influencing its decision to align with neighbouring spins.
@@
We simulate the process using the Metropolis algorithm [86, 87]. Each time, a site in the lattice is chosen randomly. The Metropolis algorithm assumes the following probability for a flipping action:
$p(\text{flip})=\begin{cases}1&\text{if }dE^{(i)}\leq 0\\ e^{-\beta dE^{(i)}}&\text{if }dE^{(i)}>0\end{cases}$ (34)
@@
Equation (34) implies that a site will always flip its spin ($p(\text{flip})=1$) if doing so results in a lower energy state. However, if the flip leads to a higher energy state, which means that the site goes from aligning to misaligning with its neighbouring sites, the coupling strength $J$ determines the site flipping likelihood. All else being equal, a higher value of $J$ results in a lower probability of flipping or higher likelihood to remain aligned with neighbouring sites.
@@ VI.2.1 Predictive information
For a given coupling strength $J$, the corresponding one-step predictive information is the mutual information between the pre-action sensory state and the post-action sensory state:
@@
Figure 13: Average predictive information plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Predictive information maximises when $J$ is small (weak coupling), and the average site exhibits explorative behaviour.
Figure 13 shows predictive information $\mathcal{I}$ for each coupling strength $J$. Predictive information is maximised when $J$ is low, suggesting that a system driven by predictive information would evolve into a highly exploratory state characterised by weak coupling between the sites. Given equation (8), predictive information can be decomposed into two terms:
$I(S_{t+1};S_{t})=H(S_{t+1})-H(S_{t+1}|S_{t})$ (37)
where the entropy $H(S_{t+1})$ represents the richness of future sensory states and the conditional entropy $H(S_{t+1}|S_{t})$ represents unpredictability.
In this example, sensory state $S$, which is the energy of a site, changes from $s$ to $-s$ if the site flips its spin. The trend observed in Figure 13 can be understood by examining the entropy and conditional entropy terms under two extreme scenarios of $J$:
  • When $J\rightarrow 0$: The weak coupling leads to frequent flipping of the sites, resulting in a sequence of alternating sensory states. Consequently, $H(S_{t+1})$ is high due to the constant activity, and $H(S_{t+1}|S_{t})=0$ because $S_{t+1}=-S_{t}$ when sites flip indefinitely. Thus, the predictive information is high, reflecting a combination of rich sensory diversity and yet fully predictable future states for a local site;
  • When $J\rightarrow+\infty$: The strong coupling prevents sites from flipping, leading to static sensory states. In this scenario,$H(S_{t+1}|S_{t})=0$ because $S_{t+1}=S_{t}$ if the site never flips, and $H(S_{t+1})\rightarrow 0$ as all sites align uniformly. Therefore, predictive information is low, indicating a predictable and homogeneous system with little sensory diversity.
Figure 14: Decomposition of predictive information into richness (dotted line) and unpredictability (dash line) components; predictive information is the difference between the two terms. Results obtained from the average over 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions.
We further note the inflexion point in Figure 13 just before the point $J_{c}=\frac{1}{2}\ln(1+\sqrt{2})\approx 0.4407$ which corresponds to the critical value for the canonical 2D-Ising model. As shown in Figure 14, $H(S_{t+1}|S_{t})$ first increases with $J$ and then decreases after the critical region, while $H(S_{t+1})$ continues to decrease, creating the inflexion point in the predictive information plot.
@@ VI.2.2 Empowerment
Figure 17: Average empowerment plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Empowerment optimises when $J$ is large (strong coupling), and the collective exhibits a stable structure. Such structure in positional configuration increases an average site’s ability to perceive the impact of its action on the environment.
Based on empowerment, the resulting optimal $J$ is at the higher end of the spectrum (Figure 17), where the lattice exhibits a more stable structure with most of the atoms aligned. Empowerment measures an agent’s ability to inject information into the environment via current actions and later retrieve the information via its sensors. A site’s action is most perceivable when all its neighbours align in the same direction, in which case the action of flip or no-flip leads to distinct sensory outputs ($s$ or $-s$, $s\neq 0$). Conversely, if four neighbours have an equal split between up and down spins, flipping the spin of a site does not change its sensory state. That is, the site will not be able to perceive the impact of its action. A large positive $J$ value increases the probability of an average site being at the configuration where all its neighbours have the same spin, thereby maximising the site’s empowerment by ensuring its actions produce noticeable changes to its future sensory inputs.
@@ VI.2.3 Variational free energy (active inference)
where the probability distribution is parameterised by coupling strength $J$.
The conditional entropy describes the expected discrepancy between the agent’s (the site’s) perception of the world (the entire lattice) and the actual future state of the world after performing action $a_{t}$. This expectation is computed based on its current internal model $m(\cdot)$. In this context, the agent’s internal model is the empirical conditional distribution of joint random variables ($S_{t+1}$, $W_{t+1}$), conditional on the values of actions $\{a_{t}\}$. It is constructed using time series $\{s_{t+1}\}$, $\{w_{t+1}\}$ and $\{a_{t}\}$.
The non-zero conditional entropy is the result of the agent’s limited sensory information, which only extends to its nearest neighbours, leading to a mismatch between the agent’s view of the world and the actual world. Under active inference, the agent would choose actions (flipping or not) to minimise this discrepancy. However, in the Ising model setting, the site can only adjust its sensitivity (coupling parameter $J$) to the influence of its neighbours. The coupling parameter effectively decides each site’s propensity to flip its spin. This parameter adjustment serves as a proxy for choosing the optimal action to reduce the conditional entropy.
Figure 18: Average negative free energy plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Negative free energy maximises when $J$ is large (strong coupling), where the spins are aligned at equilibrium, and the agent (a site)’s approximation of the world state distribution (based on local sensory states) closely matches the actual distribution that generates the world states (global property).
@@
This measure represents, on average, if a site acts with coupling parameter $J=j$, how much discrepancy it should expect between its internal model (which is based on the local sensory history) and the external world described by the lattice magnetisation. This quantity assesses how well the site’s perception aligns with the underlying global situation.
The negative free energy plot (Figure 18) reveals a similar optimal region for $J$ as the one produced by empowerment. Its negative value reflects how accurately the site’s model of the world (constructed using local information) represents the actual state of global magnetisation. When $J$ is large, more atoms are aligned, increasing the likelihood that an average site correctly perceives the overall spin direction. This alignment results in a higher action value within this range of $J$.
@@ VI.2.4 Thermodynamic efficiency
Figure 19: Average thermodynamic efficiency plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Efficiency optimises near the critical regime, where a significant portion of the work expended in tuning the parameter $J$ contributes to creating order within the system, that is, small changes of $J$ result in a large decrease in the system’s configuration entropy.
Thermodynamic efficiency reaches optimum when $J$ is near the critical value $J_{c}\approx 0.4407$, as shown in Figure 19. In this range, even a small increase in the control parameter $J$ results in a significant reduction in the system’s disorder. Consequently, the work performed to establish order in the system achieves the highest efficiency. This observation indicates that the collective systems that optimise thermodynamic efficiency at the same time operate at the critical regime.
@@ VII Principle of Super-efficiency
VIII Conclusion
This paper explored the intrinsic utility for self-organising collective systems to operate at the critical regime. We have discussed various information theoretic approaches as well as an approach that combines information theory and thermodynamics as candidates for understanding the “usefulness” of a dynamical system to operate at the critical point. An experiment was designed connecting the canonical 2D Ising model to the perception-action loop, enabling a comparison of various intrinsic utility approaches within a single example. The optimal coupling strength $J$ was computed for different approaches, including predictive information maximisation, empowerment maximisation, free energy minimisation, and thermodynamic efficiency maximisation. Each approach exhibited a distinct optimal range of parameter values, offering intuitive insights into the underlying driver shaping collective behaviour:
  • Predictive information maximises at low coupling strength, balancing sensory richness with predictability;
  • Empowerment maximises at high coupling strength, where the individuals have maximal influence over the environment;
  • Free energy minimisation (with intrinsic component only) also leads to high coupling strength, where local observations align most closely with the global configuration;

### v2

@@ I Introduction
This principle encapsulates the intrinsic utility of self-organising collective behaviour, elucidating why some systems gravitate towards criticality. Our discussion will begin with a background on criticality and phase transitions provided in Section II. Section III outlines the notations and essential technical preliminaries. Section IV provides an overview of established intrinsic utility approaches, while Section V describes thermodynamic efficiency defined at the intersection of information theory and thermodynamics. A common example using the two-dimensional Ising model is presented in Section VI to compare the self-organising behaviours driven by different intrinsic utilities. Section VII offers a more in-depth discussion on the principle of super-efficiency, and Section VIII summarises the findings and presents the final discussion. This study offers insights into the distinct characteristics of collective behaviour derived within each framework, and emphasises how the principle of super-efficiency captures the intrinsic utility for collective behaviour at the critical regime.
@@ IV.2.1 Predictive information
$\begin{split}\mathcal{I}:&=\quad I(S_{past};S_{future})\\ &=\underbrace{H(S_{future})}_{\text{Diversity}}-\underbrace{H(S_{future}|S_{past})}_{\text{Unpredictability}}\end{split}$ (8)
@@ IV.2.2 Empowerment
Equation (15) also implies dependence on the current state of the world $w_{t}$, as the context of the empowerment ($w_{t+1}$) is a combined result of the current world state $w_{t}$ and the evaluated action $a_{t}$.
@@ IV.3 Beyond intrinsic motivation
Figure 9: The diagram illustrates the interaction between elements in the active inference framework. Solid lines represent influences between components. Dash lines represent directed influence from sensory to internal or from action to external, which correspond to the two stages of active inference. Figure is adapted from [49].
@@
However, the quantity called in this approach “free energy” is different from the thermodynamic free energy. In active inference, it is instead the variational free energy formulated in terms of information-theoretic quantities, relating to the Bayesian inference process [45, 49]. Informally, anything that can be represented in the form:
@@
While equation (18) implies a connection between minimising free energy and reducing energy cost under Landauer’s principle [78], the relationship is not explicitly formulated as a ratio of informational gain to energetic costs.
@@ V Thermodynamic efficiency
At this stage, we point out that the three information-theoretic approaches reviewed in the previous section do not explicitly account for the corresponding energy costs. Thermodynamic efficiency, on the other hand, takes into consideration both the benefits and the associated costs of maintaining order within the system. Before presenting a formal definition of thermodynamic efficiency, it is important to differentiate between thermal and thermodynamic efficiency.
@@ V.1 Thermal vs thermodynamic efficiency
Let us consider a system undergoing a non-ideal process in which it receives energy and performs useful work. Not all the received energy is converted into work; some is inevitably lost as heat, which does not contribute to work output (Figure 10). Thermal efficiency measures the system’s efficiency of converting energy to work and is defined as the ratio of useful work output to total energy input, both measured in joules, rendering it a dimensionless quantity. In a non-ideal process, the second law of thermodynamics implies that this ratio is less than one.
In contrast, thermodynamic efficiency assesses the conversion of work into the system order, measured during a quasi-static change in the underlying control parameter. It pertains to systems involving interactions among multiple components and considers the benefit of increasing order within a collective system against the thermodynamic cost incurred. A system may transition from a disordered to an ordered state by altering a control parameter according to a specific protocol. Thermodynamic efficiency evaluates how efficiently the system converts the carried out work into order, at each specific value of the control parameter (Figure 11). It is quantified as the ratio of the reduction in the system’s configuration entropy (predictability gain) to the generalised work performed during the control parameter adjustment, expressed in units of bits per joule (subject to the unit of the Boltzmann constant $k_{B}$, e.g., see the expression for entropy, defined in the context of thermodynamics in Equation (21)):
@@
Figure 11: Thermodynamic efficiency for a system. It is defined as the ratio between the increase in order (measured in bits) and the change in the generalised work carried out to generate the order (measured in joules). This quantity has a unit of bits per joule.
@@ V.3 Fisher information
At this stage, we turn our attention to the Fisher Information, which is related to Gibbs free energy. The Fisher information measures the amount of information that an observable random variable carries about an unknown parameter $\theta$ which may be influencing the probability of observations. In order words, Fisher information quantifies the sensitivity of observations to the change of parameter $\theta$. Mathematically, Fisher information is defined as the variance of the score function, where the score is the derivative of the log-likelihood function with respect to $\theta$ [61]:
@@ VI.1 The 2D Ising model
The 2D Ising model offers a simplified representation of ferromagnetism in statistical mechanics. It models a collection of sites that can each exhibit either an up-spin or down-spin configuration while interacting with their neighbours to create a complex aggregate dynamic. The 2D Ising model is particularly relevant to our comparison due to its characteristic phase transition in the collective dynamics.
@@
In this example, the dynamics of the Ising model are interpreted from a perception-action loop perspective: each site acts as an agent that senses the energy of its neighbourhood and “decides” whether to flip its spin or maintain its current state. The agency of each site is determined by the choice of the coupling parameter $J$. With a high $J$ value a site is more prone to align with its neighbours, and vice versa. This parameter governs the strength of the site’s response to the neighbourhood’s energy landscape, influencing its decision to align with neighbouring spins.
@@
We simulate the process for both Gluaber dynamics [86] and Metropolis [87, 88] dynamics. At each time step, a site was selected uniformly at random. Under Glauber dynamics, the site’s spin flips with the probability:
$p_{G}(\text{flip})=0.5\left[1-\tanh(0.5\beta dE^{(i)})\right]$ (34)
Alternatively, using Metropolis dynamics the probability to flip is:
$p_{M}(\text{flip})=\min\left[1,e^{-\beta dE^{(i)}}\right]$ (35)
@@ VI.2.1 Predictive information
For a given coupling strength $J$, the corresponding one-step predictive information is the mutual information between the pre-action sensory state ($S_{t}$) and post-action sensory state ($S_{t+1}$):
@@
(a) Glauber dynamics
(b) Metropolis dynamics
Figure 13: Average predictive information plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Predictive information maximises when $J$ is small (weak coupling), where an average site exhibits explorative behaviour.
Figure 13 shows predictive information $\mathcal{I}$ for each coupling strength $J$, with two main observations highlighted. Firstly, predictive information is higher under weak coupling and decreases to nearly zero under strong coupling; this observation applies to both Glauber and Metropolis dynamics. Recall that predictive information can be decomposed into two terms:
$I(S_{t+1};S_{t})=\underbrace{H(S_{future})}_{\text{Diversity}}-\underbrace{H(S_{future}|S_{past})}_{\text{Unpredictability}}$ (38)
The overall trend is driven by the diversity component of the equation. Weak coupling promotes more exploratory behaviours, leading to diverse sensory states. In contrast, strong coupling (as $J\rightarrow+\infty$) prevents sites from flipping after settling into the lower energy state, resulting in a predictable system with little sensory diversity.
Secondly, predictive information optimises at different coupling strengths for Glauber and Metropolis dynamics – a phenomenon driven by the unpredictability term (Figure 14). Due to different behaviours of the conditional entropies in the sub-critical regime, predictive information maximises close to the critical point in Glauber dynamics, while peaks as $J\rightarrow 0$ under Metropolis dynamics. A detailed comparison of the two dynamics is provided in Appendix A.
(a) Glauber dynamics
(b) Metropolis dynamics
Figure 14: Decomposition of predictive information into richness (dotted line) and unpredictability (dash line) components; predictive information is the difference between the two curves. Results obtained from the average over 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions.
@@ VI.2.2 Empowerment
(a) Glauber dynamics
(b) Metropolis dynamics
Figure 17: Average empowerment plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Empowerment optimises at strong coupling, where the collective stabilises to a uniformly aligned configuration at equilibrium. Such configuration maximises an average site’s ability to retrieve the impact of its action through future sensory states.
Empowerment optimises at a strong coupling (Figure 17), where the lattice stabilises with most atoms aligned at the equilibrium. Empowerment measures an agent’s ability to inject information into the environment via current actions and later retrieve the information via its sensors. A site’s action is most perceivable when all its neighbours align in the same direction, in which case the action of flip or no-flip leads to distinct sensory outputs ($s$ or $-s$, $s\neq 0$). Conversely, if four neighbours have an equal split between up and down spins, flipping the spin of a site does not change its sensory state. That is, the site will not be able to perceive the impact of its action. A large positive $J$ value increases the probability of an average site being at the configuration where all its neighbours have the same spin, thereby maximising the site’s empowerment by ensuring its actions produce noticeable changes to its future sensory inputs. Empowerment is not affected by the choice of spin-flip dynamics (Glauber or Metropolis).
@@ VI.2.3 Variational free energy (active inference)
where the probability distribution $p(.)$ is parameterised by coupling strength $J$.
(a) Glauber dynamics
(b) Metropolis dynamics
Figure 18: Average negative free energy plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Negative free energy maximises at strong coupling, where the spins are aligned at equilibrium, and the site experiences minimum surprise comparing its approximate distribution of the world state (based on local sensory information) to the actual distribution that generates the world states (global property).
@@
For a given coupling strength $J=j$, this measure represents how much discrepancy an average site should expect between its internal model (based on the local sensory history) and the external world (described by the lattice magnetisation); in other words, how well the site’s local information aligns with the underlying global situation.
The negative free energy plot (Figure 18) reveals a similar optimal region for $J$ as empowerment. Maximising negative free energy effectively minimises surprise. When $J\rightarrow 0$, all the sites are actively flipping their spins. A chosen site’s local sensory states can take on all possible values, while the global magnetisation averaged to zero. This mismatch results in large surprise for an average site. On the other hand, more atoms are aligned at large $J$, increasing the likelihood that an average site correctly predicts the overall spin direction, hence reducing surprise. The same trend holds for both Glauber and Metropolis dynamics.
@@ VI.2.4 Thermodynamic efficiency
(a) Glauber dynamics
(b) Metropolis dynamics
Figure 19: Average thermodynamic efficiency plotted against different values of $J$, computed from 20 simulations, each of 20 million time steps on a $50\times 50$ square lattice with periodic boundary conditions. Thermodynamic efficiency optimises at the critical regime, where a significant portion of the work expended in tuning the parameter $J$ contributes to reducing configurational entropy; that is, the collective is most energetically efficient in creating internal order.
Thermodynamic efficiency reaches optimum when $J$ is near the critical value $J_{c}\approx 0.4407$, as shown in Figure 19. At the vicinity of the critical point, even a small increase in the control parameter $J$ results in a significant reduction in the system’s disorder. Consequently, the work performed to establish order in the system achieves the highest efficiency. This observation indicates that the collective systems that optimise thermodynamic efficiency at the same time operate at the critical regime. The same result holds for both Glauber and Metropolis dynamics.
@@ VII Principle of Super-efficiency
VIII Discussion
To summarise, this paper has explored various intrinsic utilities to understand why self-organising collective systems operate at the critical regime. We designed an experiment connecting the canonical 2D Ising model to the perception-action loop, enabling a comparison of different intrinsic utilities within a single example.
The connection is established by conceptualising each site in the Ising lattice as an agent possessing sensory-motor capabilities, thereby linking the model to the perception-action loop framework. The choice of spin-flip dynamic is analogous to the embodiment of the agents. Optimisation of the control parameter – the coupling strength $J$ – may be considered analogous to choosing the sensory channels [95] that maximises specific intrinsic utility given the embodiment.
Optimal $J$ values were computed for different approaches, including predictive information maximisation, empowerment maximisation, free energy minimisation, and thermodynamic efficiency maximisation. Each approach exhibited a distinct optimal range of parameter values, offering intuitive insights into the underlying driver shaping collective behaviour:
  • Predictive information maximises at sub-critical coupling strength, balancing sensory richness with predictability;
  • Empowerment maximises at super-critical coupling strength, where the individuals have maximal influence over the environment;
  • Free energy minimisation (with intrinsic component only) also leads to super-critical coupling strength, where local observations align most closely with the global configuration hence surprise minimised;